OpenStudy (anonymous):

find the first, fourth and eighth terms of the sequence. A(n)=-5 x 2^n-1. HELP PLEASE

OpenStudy (anonymous):

-10;-80; -1280? -5; -40 ; -640? or anything else?

OpenStudy (anonymous):

i swear i did this problem earlier today

OpenStudy (anonymous):

something similar :P

OpenStudy (anonymous):

no it was this one exactly

OpenStudy (anonymous):

hrmmm u sure? i probably never got the answer quite right. lol

OpenStudy (anonymous):

put \(n=1\) get \[-5\times 2^{1-1}=-5\times 2^0=-5\] for the first one

OpenStudy (anonymous):

second one is \[-5\times 2^{4-1}=-5\times 2^3=-5\times 8=-40\]

OpenStudy (anonymous):

ur an awesome man. Thank you.

OpenStudy (anonymous):

yw